# LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference

### v2

@@ A.3 Training datasets
Following the settings of previous LLM-based retrievers [71, 5, 47], our work uses 23 training sets pre-built by tDRO [47] or Sentence Transformers Training Data [59], including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 5 as follows.
@@
Amazon Review (2018) [53] English Amazon 999999 0.01 0.58%
cMedQA2 [81] Chinese Chinese Medical 99936 1 5.84%
DuReader [57] Chinese Chinese Web Collections 86366 1 5.05%
Eli5 [19] English Reddit 325390 0.05 0.95%
Fever [66] English Wikipedia QA 109808 0.8 5.13%
FiQA [48] English Financial 5498 1 0.32%
GooQA Pairs [30] English Web Collections 3012347 0.025 4.40%
HotpotQA [80] English Wikipedia QA 85000 0.5 2.48%
MSMARCO [52] English Web Collections 502854 1 29.39%
NFCorpus [7] English Medical 2585 1 0.15%
NPR [44] English News 594376 0.05 1.74%
NQ [36] English Wikipedia QA 58800 1 3.44%
PQA Pairs [40] English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples [25] English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract [42] English Semantic Scholar 99998 0.5 2.92%
SciFact [69] English S2ORC 806 1 0.05%
SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
StackExchange Duplicates (Title-Body) [16] English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
T2Ranking [77] Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia [28] English Wikipedia QA 60370 0.5 1.76%
Wikihow [33] English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) [82] English Yahoo 1198018 0.05 3.50%
Many downstream retrieval benchmarks [65, 76] are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 5, we slightly tune the epoch used in the dataset sampling, by following previous works in [54, 71, 47].
@@
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO [47], we perform strict training set decontamination with SimHash [49] to ensure no training queries appear in the BeIR [65] and CMTEB Retrieval [76]. The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 6.
@@
A.4 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers [71, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [55]. Following Mistral-E5 [71], we reuse most of its task instructions, and format the instruction and query in the following format:
@@ A.4 Task instructions
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [71] and tDRO [47] for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
@@
A.5 Efficiency optimizations techniques
@@
Figure 6: An example of the customized attention mask.
The customized causal mask ensures that the token blocks can attend to the common prompt area, but not attend to each other. Let the sequence length be $L$, prompt length be $P$, micro block (each token + <eos>) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index. The attention mask (0 means attend, $\infty$- ∞ means masked) is defined as,
$\text{Mask}[q,k]=\begin{cases}0,&\text{if }q<P\text{ and }k\leq q\quad\text{(attn within prompt)}\\ 0,&\text{if }q\geq P\text{ and }k<P\quad\text{(blocks can attend to prompt)}\\ 0,&\text{if }q\geq P\text{ and }\left\lfloor\frac{q-P}{w}\right\rfloor=\left\lfloor\frac{k-P}{w}\right\rfloor\text{ and }k\leq q\quad\text{(attn within current block)}\\ -\infty,&\text{otherwise}\end{cases}$
@@
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b [67] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
@@
28:         $\nabla L_{t}\leftarrow\nabla Y\odot M_{t}\in\mathbb{R}^{B\times V}$triangleright Only keep grads where $t$ was max
@@
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations  [^7] Please refer to [47]’s implementation on https://github.com/ma787639046/tdro. released by tDRO [47]. These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [41]; 2) Sequence packing [34] with Flash Attention 2 [17], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [23], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [22] for optionally enlarging batch sizes if necassary.
A.6 Detailed main results
@@
Baseline 58.0 44.0 41.2 43.2 89.2 45.5 69.3 42.7
Light Dense 49.2 38.1 37.9 34.2 83.6 40.4 60.8 39.2
Light Sparse 49.5 34.5 27.5 35.9 84.7 35.7 65.5 37.2
Light Hybrid 54.1 39.7 39.3 39.5 87.3 42.1 68.2 41.1
@@
Baseline 58.5 46.8 43.7 46.4 89.9 53.4 73.7 44.3
Light Dense 55.3 40.3 41.7 36.5 84.0 46.2 64.3 41.6
Light Sparse 46.5 34.1 26.9 38.5 86.2 38.9 66.3 39.2
Light Hybrid 56.3 41.4 42.4 41.6 87.6 46.9 69.7 42.9
@@
Baseline 57.1 49.2 44.1 48.2 90.9 59.0 77.8 46.2
Light Dense 55.8 40.9 41.5 38.7 85.7 49.8 67.6 43.4
Light Sparse 46.4 33.6 32.2 38.8 86.1 41.9 67.9 41.0
Light Hybrid 57.2 41.9 42.4 42.0 87.8 50.7 71.5 44.5
@@
Baseline 56.2 49.7 43.6 48.4 91.3 59.6 77.7 46.6
Light Dense 53.5 41.6 40.9 38.9 85.6 50.9 66.9 44.2
Light Sparse 44.0 33.6 21.9 37.0 81.0 43.6 65.4 40.9
Light Hybrid 54.3 42.2 41.4 40.5 86.9 51.7 69.9 44.8
@@
Baseline 56.6 44.1 40.4 43.9 88.9 47.9 68.1 42.6
Light Dense 51.2 38.3 36.1 33.0 82.1 41.2 59.0 39.1
Light Sparse 46.5 35.0 25.7 36.3 85.1 35.3 63.3 37.3
Light Hybrid 54.1 40.1 37.7 39.5 87.0 42.1 66.5 40.8
@@
Baseline 55.5 47.0 44.1 45.8 89.6 52.9 71.6 43.7
Light Dense 49.8 39.6 41.1 35.0 83.9 45.4 62.8 41.0
Light Sparse 45.5 36.2 28.7 36.9 84.9 38.6 64.2 38.7
Light Hybrid 52.8 41.4 42.1 39.8 86.9 46.1 68.2 42.3
@@
Baseline 57.1 48.2 43.5 47.9 90.7 56.5 74.7 44.8
Light Dense 53.3 40.0 39.1 36.8 85.1 48.8 64.4 41.8
Light Sparse 47.5 35.6 28.2 36.7 85.4 40.2 65.2 40.1
Light Hybrid 54.4 41.6 42.2 41.0 87.9 49.7 69.4 43.3
@@
Baseline 35.5 58.6 89.4 19.8 72.3 66.7 20.9 53.1
Light Dense 30.1 52.2 87.0 17.2 66.5 68.8 24.7 48.7
Light Sparse 34.0 52.0 82.2 17.2 69.1 60.6 27.8 47.6
Light Hybrid 34.2 56.6 87.4 18.6 71.5 70.5 30.2 52.0
@@
Baseline 37.6 62.8 89.5 22.5 75.3 68.7 21.2 55.6
Light Dense 31.7 56.4 87.8 19.0 70.4 68.6 21.6 51.0
Light Sparse 33.3 55.3 83.2 17.7 67.3 64.1 28.5 48.4
Light Hybrid 34.2 59.8 88.1 20.0 71.4 70.9 29.5 53.5
@@
Baseline 38.6 65.5 89.9 23.8 77.3 62.0 22.2 56.8
Light Dense 31.8 59.0 88.5 19.3 70.6 65.7 21.1 52.0
Light Sparse 33.6 58.4 84.4 18.9 69.7 63.0 27.1 49.5
Light Hybrid 34.6 61.9 88.6 20.7 72.6 71.5 27.8 54.4
@@
Baseline 36.8 66.9 90.0 22.7 77.7 71.3 23.6 57.5
Light Dense 31.9 59.8 88.7 18.5 70.3 71.5 26.7 52.7
Light Sparse 33.1 57.9 84.6 18.1 66.7 61.1 31.7 48.0
Light Hybrid 33.8 62.5 88.8 19.8 71.6 76.7 32.4 54.5
@@
Baseline 36.4 57.1 89.0 20.4 72.7 73.0 28.4 54.0
Light Dense 29.8 50.9 86.7 18.1 66.5 72.6 29.1 48.9
Light Sparse 32.9 51.6 81.3 17.4 66.4 64.8 31.1 47.3
Light Hybrid 34.0 55.8 87.0 19.3 69.2 76.8 32.3 52.1
@@
Baseline 35.9 61.4 89.5 22.3 72.9 65.7 26.0 54.9
Light Dense 30.6 54.5 87.5 18.5 64.8 71.2 26.1 50.1
Light Sparse 33.2 54.2 82.2 17.6 67.6 57.0 26.4 47.5
Light Hybrid 34.2 58.5 87.8 19.5 69.1 74.2 29.8 52.8
@@
Baseline 37.2 63.7 89.7 23.7 77.2 69.1 25.3 56.6
Light Dense 30.7 55.9 87.6 19.5 67.9 64.1 24.1 50.6
Light Sparse 33.9 55.5 78.7 18.3 69.0 60.5 28.5 48.2
Light Hybrid 34.4 60.1 87.9 21.1 71.1 73.1 29.4 53.8
@@
Baseline 32.3 74.2 84.6 53.8 73.5 50.5 78.0 63.1 63.8
Light Dense 27.2 67.6 80.8 51.0 67.9 43.0 73.4 58.5 58.7
Light Sparse 17.3 64.3 74.4 51.0 63.3 39.4 69.3 52.2 53.9
Light Hybrid 27.2 69.7 81.8 55.5 69.8 44.1 75.9 62.3 60.8
@@
Baseline 36.8 75.6 87.9 57.7 75.7 53.8 81.1 60.5 66.1
Light Dense 29.8 70.2 84.0 51.7 69.2 46.5 75.7 56.4 60.4
Light Sparse 17.2 65.4 76.1 47.4 65.0 42.5 68.5 42.0 53.0
Light Hybrid 29.9 71.5 84.8 54.5 71.2 47.2 77.1 57.7 61.7
@@
Baseline 39.5 77.1 89.5 58.9 78.1 57.6 84.1 56.2 67.6
Light Dense 32.1 71.0 85.7 53.0 70.2 49.0 78.5 52.0 61.4
Light Sparse 15.6 70.4 82.9 47.3 63.9 42.2 75.1 41.3 54.8
Light Hybrid 32.1 72.8 86.6 55.3 71.6 49.7 80.1 55.8 63.0
@@
Baseline 38.5 74.1 88.6 54.7 75.6 56.6 82.3 56.5 65.9
Light Dense 30.4 64.4 83.6 47.0 64.3 46.3 74.7 48.5 57.4
Light Sparse 11.5 63.2 80.5 43.4 61.7 36.7 71.3 37.4 50.7
Light Hybrid 30.3 67.2 84.5 50.7 67.6 46.9 76.4 51.4 59.4
@@
Baseline 37.9 78.2 87.6 56.9 75.9 54.7 78.8 57.0 65.9
Light Dense 31.3 73.8 85.3 52.8 71.8 49.5 75.2 55.7 61.9
Light Sparse 14.5 66.7 78.1 49.3 61.6 39.1 63.5 47.0 52.5
Light Hybrid 31.2 74.4 86.0 56.5 73.4 50.5 76.4 61.6 63.8
@@
Baseline 39.4 80.1 89.5 61.9 77.8 58.7 84.4 63.6 69.4
Light Dense 32.7 76.0 86.8 56.7 73.2 51.6 80.9 59.5 64.7
Light Sparse 12.8 70.4 80.2 48.8 61.1 39.4 72.4 47.9 54.1
Light Hybrid 32.6 77.0 87.6 58.3 74.0 51.6 82.0 62.8 65.7
@@
Baseline 41.3 81.4 91.0 61.1 80.0 59.8 85.5 60.3 70.1
Light Dense 34.3 74.6 88.0 55.7 75.4 53.2 81.6 57.2 65.0
Light Sparse 12.5 72.5 82.4 48.3 65.2 40.4 73.3 45.2 55.0
Light Hybrid 34.0 76.7 88.6 58.0 76.2 53.9 83.0 61.4 66.5

### v3

@@ A.5 Training datasets
Following the settings of previous LLM-based retrievers Wang et al. [2024b], BehnamGhader et al. [2024], Ma et al. [2025], our work uses 23 training sets pre-built by tDRO Ma et al. [2025] or Sentence Transformers Training Data Reimers [2019], including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 6 as follows.
@@
Amazon Review (2018) Ni et al. [2019] English Amazon $999$ 0.01 0.58%
cMedQA2 Zhang et al. [2018] Chinese Chinese Medical 99936 1 5.84%
DuReader Qiu et al. [2022] Chinese Chinese Web Collections 86366 1 5.05%
Eli5 Fan et al. [2019] English Reddit 325390 0.05 0.95%
Fever Thorne et al. [2018] English Wikipedia QA 109808 0.8 5.13%
FiQA Maia et al. [2018] English Financial 5498 1 0.32%
GooQA Pairs Khashabi et al. [2021] English Web Collections 3012347 0.025 4.40%
HotpotQA Yang et al. [2018b] English Wikipedia QA 85000 0.5 2.48%
MSMARCO Nguyen et al. [2016] English Web Collections 502854 1 29.39%
NFCorpus Boteva et al. [2016] English Medical 2585 1 0.15%
NPR Lucas et al. [2023] English News 594376 0.05 1.74%
NQ Kwiatkowski et al. [2019] English Wikipedia QA 58800 1 3.44%
PQA Pairs Lewis et al. [2021] English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples Iyer et al. [2012] English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract Lo et al. [2020] English Semantic Scholar 99998 0.5 2.92%
SciFact Wadden et al. [2020] English S2ORC 806 1 0.05%
SPECTER Cohan et al. [2020] English Semantic Scholar 136642 0.25 2.00%
StackExchange Duplicates (Title-Body) da Silva et al. [2018] English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic BeastyZ [2024] English GPT Synthetic Data 224791 1 13.14%
T2Ranking Xie et al. [2023] Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia Joshi et al. [2017] English Wikipedia QA 60370 0.5 1.76%
Wikihow Koupaee and Wang [2018] English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) Zhang et al. [2015] English Yahoo 1198018 0.05 3.50%
Many downstream retrieval benchmarks Thakur et al. [2021], Xiao et al. [2023] are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in Ni et al. [2022], Wang et al. [2024b], Ma et al. [2025].
@@
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO Ma et al. [2025], we perform strict training set decontamination with SimHash Manku et al. [2007] to ensure no training queries appear in the BeIR Thakur et al. [2021] and CMTEB Retrieval Xiao et al. [2023]. The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 7.
@@
A.6 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers Wang et al. [2024b], BehnamGhader et al. [2024] introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation Ouyang et al. [2022]. Following Mistral-E5 Wang et al. [2024b], we reuse most of its task instructions, and format the instruction and query in the following format:
@@ A.6 Task instructions
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral Wang et al. [2024b] and tDRO Ma et al. [2025] for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
@@
A.7 Efficiency optimizations techniques
@@
$\text{Input IDs}=\texttt{<bos>}~\text{Inst}~t_{0}~\texttt{<eos>}~\dots~t_{n-1}~\texttt{<eos>}$ (10)
Figure 7: An example of the customized attention mask.
The customized causal mask ensures that the token blocks can attend to the common prompt area, but not attend to each other. Let the sequence length be $L$, prompt length be $P$, micro block (each token + <eos>) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index. The attention mask (0 means attend, -$\infty$ means masked) is defined as,
@@
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b Touvron et al. [2023] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
@@
28:   $\nabla L_{t}\leftarrow\nabla Y\odot M_{t}\in\mathbb{R}^{B\times V}$ $\triangleright$ Only keep grads where $t$ was max
@@
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines BehnamGhader et al. [2024], which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations  [^4] Please refer to Ma et al. [2025]’s implementation on https://github.com/ma787639046/tdro. released by tDRO Ma et al. [2025]. These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC Li et al. [2023]; 2) Sequence packing Kundu et al. [2024] with Flash Attention 2 Dao [2023], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel Hsu et al. [2024], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache Gao et al. [2021] for optionally enlarging batch sizes if necassary.
A.8 Detailed main results
@@
Full Symmetric 58.0 44.0 41.2 43.2 89.2 45.5 69.3 42.7
LightRetriever Dense 49.2 38.1 37.9 34.2 83.6 40.4 60.8 39.2
LightRetriever Sparse 49.5 34.5 27.5 35.9 84.7 35.7 65.5 37.2
LightRetriever Hybrid 54.1 39.7 39.3 39.5 87.3 42.1 68.2 41.1
@@
Full Symmetric 58.5 46.8 43.7 46.4 89.9 53.4 73.7 44.3
LightRetriever Dense 55.3 40.3 41.7 36.5 84.0 46.2 64.3 41.6
LightRetriever Sparse 46.5 34.1 26.9 38.5 86.2 38.9 66.3 39.2
LightRetriever Hybrid 56.3 41.4 42.4 41.6 87.6 46.9 69.7 42.9
@@
Full Symmetric 57.1 49.2 44.1 48.2 90.9 59.0 77.8 46.2
LightRetriever Dense 55.8 40.9 41.5 38.7 85.7 49.8 67.6 43.4
LightRetriever Sparse 46.4 33.6 32.2 38.8 86.1 41.9 67.9 41.0
LightRetriever Hybrid 57.2 41.9 42.4 42.0 87.8 50.7 71.5 44.5
@@
Full Symmetric 56.2 49.7 43.6 48.4 91.3 59.6 77.7 46.6
LightRetriever Dense 53.5 41.6 40.9 38.9 85.6 50.9 66.9 44.2
LightRetriever Sparse 44.0 33.6 21.9 37.0 81.0 43.6 65.4 40.9
LightRetriever Hybrid 54.3 42.2 41.4 40.5 86.9 51.7 69.9 44.8
@@
Full Symmetric 56.6 44.1 40.4 43.9 88.9 47.9 68.1 42.6
LightRetriever Dense 51.2 38.3 36.1 33.0 82.1 41.2 59.0 39.1
LightRetriever Sparse 46.5 35.0 25.7 36.3 85.1 35.3 63.3 37.3
LightRetriever Hybrid 54.1 40.1 37.7 39.5 87.0 42.1 66.5 40.8
@@
Full Symmetric 55.5 47.0 44.1 45.8 89.6 52.9 71.6 43.7
LightRetriever Dense 49.8 39.6 41.1 35.0 83.9 45.4 62.8 41.0
LightRetriever Sparse 45.5 36.2 28.7 36.9 84.9 38.6 64.2 38.7
LightRetriever Hybrid 52.8 41.4 42.1 39.8 86.9 46.1 68.2 42.3
@@
Full Symmetric 57.1 48.2 43.5 47.9 90.7 56.5 74.7 44.8
LightRetriever Dense 53.3 40.0 39.1 36.8 85.1 48.8 64.4 41.8
LightRetriever Sparse 47.5 35.6 28.2 36.7 85.4 40.2 65.2 40.1
LightRetriever Hybrid 54.4 41.6 42.2 41.0 87.9 49.7 69.4 43.3
@@
Full Symmetric 35.5 58.6 89.4 19.8 72.3 66.7 20.9 53.1
LightRetriever Dense 30.1 52.2 87.0 17.2 66.5 68.8 24.7 48.7
LightRetriever Sparse 34.0 52.0 82.2 17.2 69.1 60.6 27.8 47.6
LightRetriever Hybrid 34.2 56.6 87.4 18.6 71.5 70.5 30.2 52.0
@@
Full Symmetric 37.6 62.8 89.5 22.5 75.3 68.7 21.2 55.6
LightRetriever Dense 31.7 56.4 87.8 19.0 70.4 68.6 21.6 51.0
LightRetriever Sparse 33.3 55.3 83.2 17.7 67.3 64.1 28.5 48.4
LightRetriever Hybrid 34.2 59.8 88.1 20.0 71.4 70.9 29.5 53.5
@@
Full Symmetric 38.6 65.5 89.9 23.8 77.3 62.0 22.2 56.8
LightRetriever Dense 31.8 59.0 88.5 19.3 70.6 65.7 21.1 52.0
LightRetriever Sparse 33.6 58.4 84.4 18.9 69.7 63.0 27.1 49.5
LightRetriever Hybrid 34.6 61.9 88.6 20.7 72.6 71.5 27.8 54.4
@@
Full Symmetric 36.8 66.9 90.0 22.7 77.7 71.3 23.6 57.5
LightRetriever Dense 31.9 59.8 88.7 18.5 70.3 71.5 26.7 52.7
LightRetriever Sparse 33.1 57.9 84.6 18.1 66.7 61.1 31.7 48.0
LightRetriever Hybrid 33.8 62.5 88.8 19.8 71.6 76.7 32.4 54.5
@@
Full Symmetric 36.4 57.1 89.0 20.4 72.7 73.0 28.4 54.0
LightRetriever Dense 29.8 50.9 86.7 18.1 66.5 72.6 29.1 48.9
LightRetriever Sparse 32.9 51.6 81.3 17.4 66.4 64.8 31.1 47.3
LightRetriever Hybrid 34.0 55.8 87.0 19.3 69.2 76.8 32.3 52.1
@@
Full Symmetric 35.9 61.4 89.5 22.3 72.9 65.7 26.0 54.9
LightRetriever Dense 30.6 54.5 87.5 18.5 64.8 71.2 26.1 50.1
LightRetriever Sparse 33.2 54.2 82.2 17.6 67.6 57.0 26.4 47.5
LightRetriever Hybrid 34.2 58.5 87.8 19.5 69.1 74.2 29.8 52.8
@@
Full Symmetric 37.2 63.7 89.7 23.7 77.2 69.1 25.3 56.6
LightRetriever Dense 30.7 55.9 87.6 19.5 67.9 64.1 24.1 50.6
LightRetriever Sparse 33.9 55.5 78.7 18.3 69.0 60.5 28.5 48.2
LightRetriever Hybrid 34.4 60.1 87.9 21.1 71.1 73.1 29.4 53.8
@@
Full Symmetric 32.3 74.2 84.6 53.8 73.5 50.5 78.0 63.1 63.8
LightRetriever Dense 27.2 67.6 80.8 51.0 67.9 43.0 73.4 58.5 58.7
LightRetriever Sparse 17.3 64.3 74.4 51.0 63.3 39.4 69.3 52.2 53.9
LightRetriever Hybrid 27.2 69.7 81.8 55.5 69.8 44.1 75.9 62.3 60.8
@@
Full Symmetric 36.8 75.6 87.9 57.7 75.7 53.8 81.1 60.5 66.1
LightRetriever Dense 29.8 70.2 84.0 51.7 69.2 46.5 75.7 56.4 60.4
LightRetriever Sparse 17.2 65.4 76.1 47.4 65.0 42.5 68.5 42.0 53.0
LightRetriever Hybrid 29.9 71.5 84.8 54.5 71.2 47.2 77.1 57.7 61.7
@@
Full Symmetric 39.5 77.1 89.5 58.9 78.1 57.6 84.1 56.2 67.6
LightRetriever Dense 32.1 71.0 85.7 53.0 70.2 49.0 78.5 52.0 61.4
LightRetriever Sparse 15.6 70.4 82.9 47.3 63.9 42.2 75.1 41.3 54.8
LightRetriever Hybrid 32.1 72.8 86.6 55.3 71.6 49.7 80.1 55.8 63.0
@@
Full Symmetric 38.5 74.1 88.6 54.7 75.6 56.6 82.3 56.5 65.9
LightRetriever Dense 30.4 64.4 83.6 47.0 64.3 46.3 74.7 48.5 57.4
LightRetriever Sparse 11.5 63.2 80.5 43.4 61.7 36.7 71.3 37.4 50.7
LightRetriever Hybrid 30.3 67.2 84.5 50.7 67.6 46.9 76.4 51.4 59.4
@@
Full Symmetric 37.9 78.2 87.6 56.9 75.9 54.7 78.8 57.0 65.9
LightRetriever Dense 31.3 73.8 85.3 52.8 71.8 49.5 75.2 55.7 61.9
LightRetriever Sparse 14.5 66.7 78.1 49.3 61.6 39.1 63.5 47.0 52.5
LightRetriever Hybrid 31.2 74.4 86.0 56.5 73.4 50.5 76.4 61.6 63.8
@@
Full Symmetric 39.4 80.1 89.5 61.9 77.8 58.7 84.4 63.6 69.4
LightRetriever Dense 32.7 76.0 86.8 56.7 73.2 51.6 80.9 59.5 64.7
LightRetriever Sparse 12.8 70.4 80.2 48.8 61.1 39.4 72.4 47.9 54.1
LightRetriever Hybrid 32.6 77.0 87.6 58.3 74.0 51.6 82.0 62.8 65.7
@@
Full Symmetric 41.3 81.4 91.0 61.1 80.0 59.8 85.5 60.3 70.1
LightRetriever Dense 34.3 74.6 88.0 55.7 75.4 53.2 81.6 57.2 65.0
LightRetriever Sparse 12.5 72.5 82.4 48.3 65.2 40.4 73.3 45.2 55.0
LightRetriever Hybrid 34.0 76.7 88.6 58.0 76.2 53.9 83.0 61.4 66.5